Pargo Company is preparing its master budget for 2020. Relevant data pertaining to its sales, production, and direct materials budgets are as follows.

Sales. Sales for the year are expected to total 1,700,000 units. Quarterly sales are 18%, 23%, 24%, and 35%, respectively. The sales price is expected to be $41 per unit for the first three quarters and $43 per unit beginning in the fourth quarter. Sales in the first quarter of 2021 are expected to be 15% higher than the budgeted sales for the first quarter of 2020.
Production. Management desires to maintain the ending finished goods inventories at 25% of the next quarter’s budgeted sales volume.
Direct materials. Each unit requires 2 pounds of raw materials at a cost of $12 per pound. Management desires to maintain raw materials inventories at 10% of the next quarter’s production requirements. Assume the production requirements for first quarter of 2021 are 494,000 pounds.


Prepare the sales, production, and direct materials budgets by quarters for 2020.Pargo Company is preparing its master budget for 2020. Relevant data pertaining to its sales, production, and direct materialPARGO COMPANY Direct Materials Budget Quarter A IA
Answer #1

a)

PARGO COMPANY

SALES BUDGET

FOR THE YEAR ENDED DECEMBER 31 2020

QUARTER
1 2 3 4 Year
unit sales 1700000*18%= 306000 1700000*23%= 391000 1700000*24%= 408000 1700000*35%= 595000 1700000
Unit price 41 41 41 43
Total sales (a*b) 12,546,000 16,031,000 16,728,000 25,585,000 70,890,000

b)

PARGO COMPANY

DIRECT MATERIAL BUDGET

FOR THE YEAR ENDED DECEMBER 31 2020

Quarter
1 2 3 4 Year
units to be produced 327250 395250 454750 534225
material requirement per unit 2 2 2 2
Total material required for production 654500 790500 909500 1068450
Desired ending raw material inventory 790500*10%=79050 90950 106845 49400 (494000*10%)
Total needs 733550 881450 1016345 1117850
less:Beginning raw material inventory - 65450   (654500*10%) -79050 -90950 -106845
Material needs to be purchased 668100 802400 925395 1011005
cost per pound 12 12 12 12
Total cost of material purchase 8017200 9628800 11104740 12132060 40,882,800

Working:

1)Sales of first quarter of 2021 = sales of first quarter of 2020 (1+%increase)

                    = 306000 (1+.15)

                    = 351900

2)

PRODUCTION BUDGET
1 2 3 4 1 quarter of 2021
Unit sales 306000 391000 408000 595000 351900
Desired ending finished goods inventory 391000*25%= 97750 102000 148750 87975
less:Beginning finished goods inventory -76500 [306000*25%] -97750 -102000 -148750
Units to be produced 327250 395250 454750 534225